Density of water
Mass of flask and stopper + water 104.915 g
Mass of flask and stopper 87.175g
Mass of water ?
Volume of water 10.00 g

Show the calculation for the density of water for trial 1

Density of water? g/mL
1
Expert's answer
2016-09-16T09:04:03-0400
Mass of water = 104.915 g - 87.175g = 17.74 g
If the volume of water is 10 mL and mass 17.74 g we can calculate the density of water:
d(H2O) = m (H2O) / V(H2O) = 17.74 g / 10 mL = 1.774 g/mL
Answer: 1.774 g/mL